Project

General

Profile

Observing Plan for MJD 56323 (starting evening of January 30 2013)

Observers

Observer 1 Observer 2 Run Manager
Old/March Kunder/Rooney Soares-Santos

DECam Configuration and Setup

SISPI Configuration File sv.ini
DECam constants DECAM:CURRENT
Proposal ID 2012B-0003
Image Processing (Calibration) NCSA
Image Processing (First Cut) NCSA

Almanac and Weather

Twilight Evening: 01:12 UT
Twilight Morning: 08:42 UT
Moon Rise: 22:32
Moon Set: 11:28
Moon Illum.: 95 %
Middle of night hand over to community at 01:57

Plan and Goals for the Night

  1. Afternoon calibration data:
    1. 5 x 60 sec dark exposures to check for any changes in the hot-pixel list.
    2. dome flats including u-band in case we get to the half-chip SNe observations.
  2. Calibration observations (Douglas Tucker):
    1. Attached (FebStds.txt) is a list of night-by-night suggestions for the rest of the DES time, assuming the normal evening twilight times and a single standard star field just before middle-of-the-night switchover.
    2. Currently, we are observing 3 standard star fields in evening twilight -- at least one and preferably more in SDSS Stripe 82 -- and one more just before switchover. For the 3 standard fields observed during evening twilight, one should be at low airmass (<1.25), one at medium airmass (1.25-1.65), and one at high airmass (1.65-2.10).
    3. BEFORE taking/queuing the evening twilight standards tonight, do a quick check by running on observer2:
      source ~/dtucker/bin/setup_StdStarFieldPicker.csh (or .bash if you are using bash shell)
      decamStdFieldsTimeDate.py --UT='2013/01/30 00:43' (replacing the time with the anticipated time that the standards sequence will start)

      and, if SDSSJ0100-0000 does not appear in the resulting table, replace it with an observation of the E2-A prenight script
  3. SN observations (Masao Sako):
    Given the shortage of observing time, the SN group agreed that it will be more useful to obtain 2 more epochs of half of our fields than to get 1 epoch of each. So our plan is to drop E1,E2,X1,X2,X3 and continue with S1,S2,C1,C2,C3. Each of the 5 fields are in the deadman state. Here are the priorities for tonight.
    1. C3 z (12 days old)
  4. Obstac (Eric Neilsen):
    If seeing is <1.3 arcsec, run in auto-mode to observe SPT-E mini-survey.
  5. Half-chip SN observations (Huan Lin):
    Exposures of each SN field taken with 1/2-chip shift from normal pointings, to reconcile the photometric systems. They can be done in poor seeing but should not have clouds. On the night these are done, you will need to take u-band dome flats and add u-band to the standards as well. These exposures are lower priority than obtaining the regular SN observations above.
    1. X3, C3 (VVDS, CDFS)
    2. X1, C1, C2 (recent AAT redshifts)
    3. X2, S1, S2, E1, E2 (only if there is time; check if not too low)

Notes from the 4PM Meeting and Additional Comments

  • please keep an eye on the guider traces and try to estimate the rate at which oscillations appear vs declination of the target. We would like to see how the horseshoe damper motor upgrade is influencing tracking issues.
  • When observing the SN fields, please record in the e-log the offsets that you apply to the pointings (via the observer's console) and the results of the "center" command. There should no longer be any need for cos(dec) corrections. Also please do not iterate more than once on the SN field pointings; this consumes a lot of time.
  • No need to fine-tune the pointings for the SN exposures that will have the half-chip shifts.
  • Keep an eye on the wind speed, especially when the dome is facing upwind (usually North). Alistair noted winds easily inducing telescope motion (again you can watch the guider traces). Basic instructions: for winds > 5-8 mph, raise the wind blind to cover ~ 30% of the telescope. For winds >15 mph, close the louver windows in the dome as well.
  • See the SV "extended SV schedule" page for instructions on how soon you can begin standards and sky observations.